OpenStudy (anonymous):

Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). (3a3 − 5b3)+ ? a3+ ? b3 = (a3 + b3).

OpenStudy (anonymous):

@ajo21

OpenStudy (anonymous):

sorrryyy I have no clue!!!!!

OpenStudy (anonymous):

@usercode3rror

OpenStudy (anonymous):

\left(3a^3\:−\:5b^3\right)+\:a^3+\:b^3=\:\left(a^3+\:b^3\right).

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

4a3−4b3

OpenStudy (anonymous):

4a3−4b3= (a3 + b3).

OpenStudy (anonymous):

I NEEED HELP

OpenStudy (anonymous):

MY CALC NOT working

OpenStudy (anonymous):

on what

OpenStudy (anonymous):

okay wait

OpenStudy (retireed):